OpenStudy (anonymous):

reduce the equation 3x + 6y -2 =0 to the slope-intercept form

OpenStudy (anonymous):

reduce the equation 3x + 6y -2 =0 to the slope-intercept form

OpenStudy (anonymous):

slope-intercept form is y = mx + c ?

OpenStudy (anonymous):

what is the solution?

OpenStudy (anonymous):

if it is y = mx + c (as google says :p) you have 3x + 6y -2 = 0 first you let y be alone on the left : 6y = 2 - 3x and then you divide by 6 : y = 2/6 - 3x/6 wich is : y = 1/3 - x/2

OpenStudy (anonymous):

i can give more details if u need to

OpenStudy (anonymous):

y = mx + c Standard slope intercept form.

OpenStudy (anonymous):

yes i need to please. how about this...what is the equation of the line that is passes through the origin and parallel to the line 5x-3x+2=0?

OpenStudy (anonymous):

Post another question for that.

OpenStudy (anonymous):

what details do you nedd on that particular question ?

OpenStudy (anonymous):

@estudier which one?

OpenStudy (anonymous):

@niwi yes

OpenStudy (anonymous):

"how about this...what is the equation of the line that is passes through the origin and parallel to the line 5x-3x+2=0? Post as separate question, thread is confusing otherwise..

OpenStudy (anonymous):

where do you need details ? tell me what do you understand or i could spend hours explaining o_o

OpenStudy (anonymous):

what you don't understand*

OpenStudy (anonymous):

@estudier i posted already